### v2

# Graph neural network-based structural classification of glass-forming liquids and its interpretation via self-attention mechanism

Abstract
Glass-forming liquids exhibit slow dynamics below their melting temperatures, maintaining an amorphous structure reminiscent of normal liquids. Distinguishing microscopic structures in the supercooled and high-temperature regimes remains a debated topic. Building on recent advances in machine learning, particularly Graph Neural Networks (GNNs), our study automatically extracts features, unveiling fundamental mechanisms driving structural changes at varying temperatures. We employ the self-attention mechanism to generate attention coefficients that quantify the importance of connections between graph nodes, providing insights into the rationale behind GNN predictions. Exploring structural changes with decreasing temperature through the GNN+self-attention using physically-defined structural descriptors, including the bond-orientational order parameter, Voronoi cell volume, and coordination number, we identify strong correlations between high attention coefficients and more disordered structures as a key indicator of variations in glass-forming liquids.
I Introduction
When a liquid is cooled below its melting temperature, it enters a supercooled state while retaining its disordered amorphous structure. As the temperature continues to decrease toward the glass transition temperature, the structural relaxation time experiences a rapid increase, eventually leading to solidification. The phenomenon is commonly referred to as the glass transition. Theoretical and computational elucidation aimed unraveling the characteristics that govern the slow dynamics in glass-forming liquids, particularly within structures closely similar to those of normal liquids, has presented a formidable problem. [1]
The challenge of unraveling the structural order governing slow dynamics in glass-forming liquids has been met with recent breakthroughs. In fact, this advancement employs machine learning (ML) techniques, enhancing predictive accuracy and unveiling fundamental mechanisms driving structural changes. [2, 3] ML prove exceptional in detecting subtle temperature-related structural changes, even with varying relaxation times, revealing a crucial link between structural features and slow dynamics in glass-forming liquids.
Specifically, ML methodologies have been applied for data obtained from molecular dynamics (MD) simulations to explore multifaceted origins of the slow dynamics from the atomistic level. [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43] These investigations have effectively opened new perspectives for comprehending the characteristics between liquids and glasses, spanning the prediction of their properties to the elucidation of structural and dynamical changes by varying temperatures. The seamless integration of ML and MD represents a transformative approach, which yields insights into the behaviors inherent to various glass-forming liquids.
More recently, there has been a notable increase in the application of sophisticated ML techniques, including Convolutional Neural Networks (CNNs) [44] and Graph Neural Networks (GNNs) [45, 46], as evidenced in literatures. In particular, the GNN is a neural network model designed to exploit the capabilities of ML within graph structure, where the information is represented in terms of nodes and edges. It iteratively updates information associated with the nodes and edges present in the input graph through message passing. Therefore, the GNN eliminates the requirement for introducing features manually, allowing automatic discovery of relevant structural information. Bapst et al. achieved remarkable success in accurately predicting the dynamic propensity in a three-dimensional glass-forming liquids from its static structure, utilizing the power of GNNs. [12] The dynamic propensity refers to the identification of single-particle displacements exhibiting higher mobility than the average. [47, 48, 49] Consequently, their work established a compelling link between the dynamics and structural attributes of glass-forming liquids. Shiba et al. also succeeded in predicting the dynamic propensity with higher accuracy than the results of Bapst et al. [25] Their success can be attributed to an extended approach that involved learning not only individual particle displacements but also the pair-particle displacements as part of the GNNs output. Additionally, other advanced GNN techniques were described, including the incorporation of a geometry-enhanced GNN [33] and a rotation-equivariant graph structure. [35] These studies demonstrated that the GNNs can automatically extract features from graph-based inputs through message passing, without the need for predefined physical descriptors. However, their interpretability remains limited due to their black-box nature.
Swanson et al. pursued an alternative approach and found that CNNs and GNNs possess the capability to discriminate between liquid and glass structures across the glass transition temperature $T_{\mathrm{g}}$ of the two-dimensional glass-forming liquids. [13] Notably, the self-attention mechanism, a method within eXplainable Artificial Intelligence (XAI), provides insights into how GNNs differentiate liquids and glasses. Originally introduced in Graph Attention Networks, the self-attention mechanism quantifies the importance of a given node in relation to its edge-connected neighbors within the graph structure, providing interpretability to black-box nature of GNN. [50, 51] Other studies have also highlighted the effectiveness of the self-attention mechanism in ML, revealing connections between structure features and dynamic heterogeneities. [33, 36, 41]
Furthermore, Oyama et al. demonstrated comparable liquid-glass classification using CNNs. [28, 39] They further clarified the discriminative features underlying CNNs’ predictions by employing the Gradient-weighted Class Activation Mapping (Grad-Cam) method. The Grad-Cam approach provides a local-explanation model for CNN predictions, [52] revealing a correlation between the spatial distribution of the Grad-Cam score and that of the dynamic propensity.
Figure 1: Schematic picture of the GNN architecture conducted in this study. The node and edge features, $\vec{h}^{\mathrm{in}}_{v}$ and $\vec{h}^{\mathrm{in}}_{e_{v,u}}$ are first generated from the graph data of the particle configuration obtained through MD simulations. The GNN architecture includes the encoder, $n$ cycles of edge and node updates, self-attention layer, decoder, and FCNN for predicting the binary classification between 0 (representing temperature $T_{1}$) and 1 (representing temperature $T_{2}$). The index $v$, $u$, $s$ refer to node indices within the range of 1 to $N$, where $N$ represents the total number of particles. In this figure, 3 nodes are written, while in the calculations actually performed, the number of nodes is $N=4096$.
It is worth noting that the aforementioned studies have indeed showcased the success of ML and XAI in learning structural features in glass-forming liquids. However, GNNs have yet to fully capture the temperature-dependent structural changes of glass-forming liquids, particularly for three-dimensional systems, or provide a comprehensive physical interpretation of the nodal features generated from GNNs.
In the present study, we employ GNNs to classify the structures of a three-dimensional glass-forming liquid model consisting of two types (type 1 and type 2) of particles of different sizes.[53, 54, 55] Our objective is to distinguish two temperature states based solely on particle configurations, along the supercooled liquid branch, which lies above the glass transition temperature. In this regard, our approach differs from previous studies by Swanson et al. and Oyama et al., which used a constant cooling rate distinguishing liquid and glass structures. [13, 28, 39]
We further explore the rationale behind GNN predictions by adopting the self-attention mechanism, with the aim of elucidating the distinctive structural changes that occur as temperature varies. Specifically, we compare node features generated by GNNs with physically-defined structural descriptors, including bond orientational orders (BOOs), [56, 57, 58, 59] Voronoi cell volume, [60] and coordination number (C.N.), to interpret how graph-based ML differentiates the structures of glass-forming liquids at different temperatures. These structural descriptors characterize the nearest neighbor particle environment, quantifying the degree of the crystalline order. These parameters are expected to serve as salient indicators governing the slow dynamics in glass-forming liquids. [61] Additionally, we compare the attention coefficients, which quantify the importance of a target node relative to its neighboring nodes, with these structural descriptors to provide a physical interpretation of the underlying mechanism in GNNs predictions.
II Methods
II.1 Molecular dynamics simulations
The model employed was a well-known model of glass-forming liquids, namely, the three-dimensional binary soft-sphere (SS) model. [53, 54, 55] In the SS model, the interaction potential is based on
$\phi(r)=\varepsilon_{\alpha\beta}\left(\frac{\sigma_{\alpha\beta}}{r}\right)^{12},$ (1)
and the truncated and force-shifted interaction, defined as:
$U(r)=\phi(r)-\phi(r_{\mathrm{c}})-(r-r_{\mathrm{c}})\left.\frac{d\phi(r)}{dr}\right|_{r=r_{\mathrm{c}}},$ (2)
is employed to ensure that both the potential and the force approach zero continuously at the cutoff length $r_{\mathrm{c}}=2.5\sigma_{\alpha\beta}$. The size ratios are defined as $\sigma_{22}/\sigma_{11}=1.2$ and $\sigma_{12}/\sigma_{11}=1.1$. Additionally, the mass ratio is represented as $m_{2}/m_{1}=2$, while the equal energy scale is used with $\varepsilon_{11}=\varepsilon_{12}=\varepsilon_{22}$. The reduced units for length, temperature, and time are $\sigma_{11}$, $\varepsilon_{11}/k_{\mathrm{B}}$, and $\sqrt{m_{1}\sigma_{11}^{2}/\varepsilon_{11}}$, respectively. The ratio of particle numbers between type 1 and type 2 is set to 50:50 within a total particle number of $N=4096$. The total number density is fixed at $\rho=0.75$, and the investigated temperatures are $T=1.0$, $0.8$, $0.6$, $0.4$, $0.3$, $0.25$, $0.24$, $0.23$, $0.22$, and $0.21$. A time step $\Delta t=0.005$ was used. All MD simulations were computed using Large-scale Atomic/Molecular Massively Parallel Simulator (LAMMPS). [62]
The results of the radial distribution function $g(r)$ and the self-part of the intermediate scattering function $F_{s}(k,t)$ are found in Supplementary Fig. S1(a) and S1(b) of the supplementary material, respectively. Furthermore, the temperature-dependent behavior of the structural $\alpha$-relaxation time $\tau_{\alpha}$, which was determined from $F_{s}(k,t)$, is illustrated in Fig. S1(c) of the supplementary material. The onset temperature $T_{0}$, below which $F_{\mathrm{s}}(k,t)$ begins to develop a two-step relaxation, was estimated to be $T_{0}=0.3$[63] Furthermore, based on the power-law divergence behavior predicted by the mode-coupling theory, the critical temperature was estimated as $T_{\mathrm{c}}=0.198$[63] Another glass transition temperature, identified as the divergence temperature of $\tau_{\alpha}$, was estimated as $T_{\mathrm{K}}=0.181$ using the Vogel–Fulcher–Tammann equation, $\tau_{\alpha}=\tau_{0}\exp(D/(T-T_{\mathrm{K}}))$, as shown in Fig. S1(c). Thus, the temperature range examined in this study involves a deeply supercooled regime above the glass transition temperature. The temperature dependence of the potential energy is displayed in Fig. S1(d) of the supplementary material.
II.2 GNN and self-attention mechanism
The schematic picture of our GNN architecture is depicted in Fig. 1. The particle configuration can be represented by embedding features of particle types and connections between pairs of particles as attributes of nodes and edges of the graph structure, respectively. Specifically, we utilize the particle type (type 1 or type 2) as input feature vectors for the graph nodes. A pair of particles are treated as neighbors, forming an edge when their distance is smaller than $r_{\mathrm{cut}}=1.8$. The relative coordinates of the pair of neighboring particles serve as input features for the edges. The value of $r_{\mathrm{cut}}=1.8$ corresponds to the first minimum length of the radial distribution functions $g(r)$ of large particles (type 2). Note that this GNN approach is particularly appropriate when applied to many particle systems since it effectively retains the three-dimensional structural information.
Consider a graph $G(V,E)$, where $V$ represents the set of nodes and $E$ denotes the set of edges. For any pair of nodes $v$ and $u$ belonging to $V$, an edge connecting $v$ and $u$ is denoted as $e_{v,u}$. The collection of nodes neighboring a node $v\in V$ is represented as $N(v)$. The initial step involves an encoding procedure using the input feature vectors $\vec{h}^{\mathrm{in}}_{v}$ and $\vec{h}^{\mathrm{in}}_{e_{v,u}}$ for a node $v\in V$ and an edge $e\in E$, respectively. This procedure is described as follows:
$\vec{h}^{0}_{v}$ $=\mathrm{EN}(\vec{h}^{\mathrm{in}}_{v}),$ (3)
$\vec{h}^{0}_{e_{v,u}}$ $=\mathrm{EN}(\vec{h}^{\mathrm{in}}_{e_{v,u}}),$ (4)
where $\mathrm{EN}(\cdot)$ represents the encoder function. In practice, $\mathrm{EN}(\cdot)$ consists of one hidden layer and the output layer, both utilizing the Rectified Linear Unit (ReLU) activation function. This encoder generates $\vec{h}^{0}_{v}$ and $\vec{h}^{0}_{e_{v,u}}$, each composed of 32 elements for node $v$ and edge $e_{v,u}$.
Secondly, features associated with nodes and edges within the graph are mutually calculated, and updates to both node and edge features undergo iteration for $n$ cycles. This procedure is described as follows:
$\vec{h}^{m}_{e_{v,u}}$ $=\mathrm{MLP}\left(h^{m-1}_{e_{v,u}}\oplus h^{m-1}_{v}\oplus h^{m-1}_{u}\right),$ (5)
$\vec{h}^{m}_{v}$ $=\mathrm{MLP}\left(\vec{h}^{m-1}_{v}\oplus\sum_{u\in N(v)}\vec{h}^{m}_{e_{v,u}}\right),$ (6)
where $\mathrm{MLP}(\cdot)$ denote a multi-layer perception function, consisting of one hidden layer and the output layer, both utilizing the ReLU activation function. Here, $m$ denotes the index of the cycle, i.e., $0<m\leq n$, and the number of cycles was set to $n=10$. This message passing procedure for updating nodes and edges propagates the information from neighboring particles, eventually encompassing the entire system.
We introduce the self-attention layer into our GNN architecture in order to assess which node’s information to emphasize, considering the information exchanged among neighboring nodes. [50, 51] Note that Swanson et al. employed the self-attention layer at the end of message passing in Ref. 13, which is compatible with our GNN architecture, as shown in Fig. 1. The attention coefficient $e_{vu}$ is defined as
$e_{vu}=a(\mathbf{W}\vec{h}^{n}_{v},\mathbf{W}\vec{h}^{n}_{u}),$ (7)
which measures the importance of node $u$’s feature to node $v$. Here, $\mathbf{W}$ is a learnable weight matrix and $a$ signifies the so-called shared attentional mechanism. In the procedure of $a$, the concatenation of two vectors, $\mathbf{W}\vec{h}^{n}_{v}$ and $\mathbf{W}\vec{h}^{n}_{u}$, is fed into to a single-layer neural network, which is parameterized by a weight vector $\vec{a}$, and further employs the Leaky ReLU nonlinearity (negative input slope of 0.2). Using the Softmax function, the attention coefficient is normalized as follows:
$\alpha_{vu}=\mathrm{Softmax}(e_{vu})=\frac{\exp(e_{vu})}{\sum_{s\in\mathcal{N}(v)}\exp(e_{vs})},$ (8)
where $\mathcal{N}(v)$ represents the neighboring nodes of $v$, including node $v$ itself. Note that $\alpha_{vu}$ is summed over $u$ to unity and is asymmetric with respect to the node exchange, i.e., $\alpha_{vu}\neq\alpha_{uv}$ for $v\neq u$, since the graph is directed. Finally, new node features
$\vec{H}_{v}=\alpha_{vv}\mathbf{W}\vec{h}^{n}_{v}+\sum_{u\in\mathcal{N}(v)}\alpha_{vu}\mathbf{W}\vec{h}^{n}_{u},$ (9)
are generated from the self-attention layer.
For each node, the output feature $\vec{H}_{v}$ with 32 elements is transformed into a single element variable $H^{\prime}_{v}$ through the decoder that incorporates two hidden layers and employing the ReLU, and an output layer with the linear transformation. The node feature $H^{\prime}_{v}$ (4096 nodes) are combined and undergo training through a fully connected neural network (FCNN). This FCNN includes a single hidden layer utilizing the ReLU.
Through the encoder, message passing, self-attention layer, decoder, and FCNN, a binary output predicts a value of 0 for the structure at $T_{1}$ and 1 at $T_{2}$ following a sigmoid function. In the context of the binary classification, the loss function employed is binary cross-entropy, which is minimized during the training process using the training dataset.
We used PyTorch for implementation, utilizing the Adam optimization algorithm. The learning rate and batch size were set to $10^{-6}$ and 128, respectively. These hyperparameters were selected to ensure sufficient optimization of the loss function, as discussed in the following paragraph. The maximum number of epochs was set to 10000, and the training was terminated when the loss function of the test dataset was minimized, at which point the GNN was applied for prediction.
The number of weights in the GNN used in this study is 4096 $\times$ 32 for node features and 4096 $\times$ 32 $\times$ (number of coordination numbers per particle, ranging from 10 to 25) for edge features. We obtained 500 equilibrium particle configurations at each temperature from MD simulations. Of these, 400 were allocated for training dataset, while the remaining 100 were used for test dataset. In total, each GNN model for a given $(T_{1},T_{2})$ combination utilized 800 configurations (400 from $T_{1}$ and 400 from $T_{2}$) used for training, and 200 configurations (100 from $T_{1}$ and 100 from $T_{2}$) for testing.
II.3 Structural descriptors for local particle environment
The BOOs are useful to characterize the local orientational orders of liquid and crystal structures. [56, 57, 58, 59] To characterize the orientational symmetry among the neighbors of particle $i$, the complex parameter is defined as
$q_{lm}(i)=\frac{1}{N_{\mathrm{b}}(i)}\sum_{j=1}^{N_{\mathrm{b}}(i)}Y_{lm}(\hat{\bm{r}}_{ij}),$ (10)
where $N_{\mathrm{b}}(i)$ represents the number of first nearest neighboring particles of particle $i$, within a cutoff radius of $r_{\mathrm{cut}}=1.8$. This length corresponds to the length scale at which neighboring nodes are connected by edges in the GNN. Furthermore, $Y_{lm}(\cdot)$ are the spherical harmonics of degree $l$ and order $m$ (running from $m=-l$ to $m=+l$). $\hat{\bm{r}}_{ij}$ represents the relative unit vector between particles $i$ and $j$. To consider the rotational invariance, the local BOO is defined as follows:
$Q_{l}(i)=\sqrt{\frac{4\pi}{2l+1}\sum_{m=-l}^{l}|q_{lm}(i)|^{2}}.$ (11)
While the selection of the order $l$ is arbitrary, $Q_{6}$ and $Q_{4}$ are often preferred due to its ability to differentiate between liquid and crystal structures. More specifically, $Q_{6}$ is sensitive to the hexagonal symmetry and tends to be large for FCC (face-centered cubic) and HCP (hexagonal close-packed) structures, while $Q_{4}$ is sensitive to structures with local cubic symmetry. We employed the Pyboo code for the computation of $Q_{6}$ and $Q_{4}$[64]
Note that the coarse-grained BOO has been proposed by averaging $q_{lm}(i)$ over neighboring particles of particle $i$ and particle $i$ itself. [65] This approach has been shown to provide a more precise and discriminative characterization between liquid and crystal structures when compared to the original BOO method. We have evaluated the use of this coarse-grained BOO in our neural network; however, we have opted not to employ it in this study. Our observations indicated that the additional averaging step tends to obscure the distinctions among disordered structural configurations within the systems we are investigating. Note that an analogous observation has been documented in a previous study, where unsupervised machine learning was employed for dimension reduction of local structural orders in glass-forming liquids. [22]
An alternative approach to characterize local orders is based on Voronoi tessellation. [60] In particular, the Voronoi cell serves as an indicator of the local geometry associated with each particle in glass-forming liquids. [66, 67] We conducted the Voronoi cell volume calculations using the Pyvoro code. [68] A final structural descriptor examined in this study is C.N. within the first coordination shell, defined as the number of particles within $r_{\mathrm{cut}}=1.8$.
The temperature dependence of the probability density distributions for $Q_{6}$, $Q_{4}$, Voronoi cell volume, and C.N. is shown in Fig. S2(a), (b), (c), and (d) of the supplementary material, respectively. Figure S2(a) and (b) reveal that $Q_{6}$ increases while $Q_{4}$ decreases as the temperature decreases, indicating the development of hexagonal crystalline order. Figure S2(c) shows that the probability distribution of the Voronoi cell volume exhibits an approximately Gaussian shape at high temperatures. However, as the temperature decreases, the distribution develops two distinct peaks, indicating increasing spatial heterogeneity in the local structure. Specifically, the distribution of small (type 1) and large (type 2) particles become distinct as the supercooled liquid approaches the glass transition. Figure S2(d) demonstrates C.N. tends to decrease slightly with decreasing the temperature. Note that, although this result may seem counterintuitive, the value is highly dependent on the choice of the cutoff distance, $r_{\mathrm{cut}}$. A similar dependence on $r_{\mathrm{cut}}$ has been observed for BOO values. [69]
Figure 2: Fraction of GNN prediction accuracy (a) and AUC (b) for classification between two temperatures, $T_{1}$ and $T_{2}$.
Figure 3: Pearson correlation coefficients between node features, denoted as $H^{\prime}_{u}$, generaged by GNN, and structural descriptors including two BOOs, $Q_{6}$ (a), $Q_{4}$ (b), Voronoi cell volume (c), and C.N. (d) for each particle. The temperature combinations are identical to those presented in Fig. 2.
Figure 4: Fraction of FCNN prediction accuracy for classification between two temperatures, $T_{1}$ and $T_{2}$, using two BOOs, Q6 (a), Q4 (b), Voronoi cell volume (c), and C.N. (d) for each particle.
Figure 5: Probability distribution plots using two BOOs, $Q_{6}$ (a) and $Q_{4}$ (b), Voronoi cell volume (c), and C.N. (d) for temperature combination of $T_{1}=0.21$ and $T_{2}=1.0$. The distributions are colored by the corresponding coordination-weighted attention coefficient $\tilde{\alpha}_{vu}$ values given in the color bar.
III Results and discussion
III.1 GNN classification of different temperatures
Figure 2(a) and (b) present the prediction accuracy and the receiver operating characteristic area under the curve (ROC-AUC, hereafter referred to simply as AUC) by our structure-based GNN temperature predictions, evaluated across various combinations of two temperatures, $T_{1}$ and $T_{2}$, respectively. In each combination of $T_{1}$ and $T_{2}$, $T_{1}$ is selected to be a lower temperature than $T_{2}$. Here, in the interpretation of prediction output, we assign the label $T_{1}$ if the output yields a value less than 0.5, and $T_{2}$ otherwise. The results demonstrate that the structure at each of the two temperature combinations examined in this study can be perfectly distinguished from one another. The evolution of the epoch-dependent loss function for both the training and test datasets is depicted in Fig. S3 of the supplementary material. Notably, the GNN successfully classifies different temperatures, even when the radial distribution function $g(r)$ shows small variations, while the $\alpha$-relaxation time $\tau_{\alpha}$ increases significantly as the temperature decreases (see Fig. S1(a), (b), and (c) of the supplementary material). However, a slight decrease in the prediction accuracy was observed for combinations of higher temperatures, e.g., $T_{1}=0.8$ and $T_{2}=1.0$. This result implies that the GNN predictions are accompanied by a specific level of resolution in the structure classification across varying temperature. To explain this aspect further, we analyzed the prediction accuracy and AUC of GNNs for supercooled states within the temperature range 0.21-0.25, as shown in Fig. S4 of the supplementary material, with the corresponding epoch-dependent loss function presented in Fig. S5 of the supplementary material. A decline in predictive performance was particularly evident for closely spaced temperature combinations, consistent with insufficient optimization of loss function. These findings suggest that closely spaced temperatures result in overlapping fluctuations in the potential energy, making it challenge for GNNs to distinguish between structures between two different temperatures (see Fig. S1(d) of the supplementary material). This limitation may be improved by using inherent structures obtained through energy minimization, which would reduce reducing energy overlap.
III.2 Node features generated by GNN
We next examined the physical properties of the node features generated by the GNN, denoted as $H^{\prime}_{u}$ for the node index $u$ in Fig. 1, which were were subsequently employed in the FCNN to predict temperature. For this aim, we calculated BOOs ($Q_{6}$ and $Q_{4}$), Voronoi cell colume, and C.N. for each particle of a specific configuration and compared these structural descriptors with the GNN-generated node features $H^{\prime}_{u}$ using the same configuration. Specifically, the Pearson correlation coefficient was calculated for each $(T_{1},T_{2})$ pair snapshot and averaged over 200 configurations (100 configurations from $T_{1}$ and 100 configurations of $T_{2}$ from the test dataset). The results are illustrated in Fig. 3. Overall, $Q_{6}$ and Voronoi cell volumes display positive correlations with the node features reaching values approximately 0.3, while C.N. exhibit negative correlations of up to 0.4 in magnitude. Note that $Q_{4}$ also shows slight positive correlations, though less procounced than those of $Q_{6}$. A more detailed observation reveals that, for all ($T_{1}$, $T_{2}$) pairs, the absolute values of the correlations with node features generated by GNNs follow the order: C.N. $>$ Voronoi cell volume $>$ $Q_{6}$ $>$ $Q_{4}$. This trend can be attributed to the methodology used to compute structural descriptors: $Q_{6}$ and $Q_{4}$ incorporate angular information of neighboring particles. In contrast, C.N. and Voronoi cell volume provide a more direct capture of local structural environment. The observed correlations between node features and structural descriptors suggest that the GNN’s success in two-temperature discrimination arises from its ability to effectively capture insights into the local particles environment.
III.3 FCNN classification of different temperatures using structural descriptors
The correlations between GNN-generated node features and physically-defined structural descriptors suggest that ML with the FCNN, utilizing these structural descriptors, may also effectively discriminate between the two temperatures. In fact, the distribution of each structural descriptor exhibits some temperature dependence, as shown in Fig. S2 of the supplementary material, suggesting that FCNN can presumably distinguish temperature differences.
We thus implemented an FCNN to classify the structure between two temperatures. The architecture consists of an input layer with 4096 nodes, a single hidden layer also with 4096 nodes utilizing the ReLU, and a binary output layer predicting the temperature ($T_{1}$ or $T_{2}$) via a sigmoid function. The model is trained by minimizing binary cross-entropy. Input variables for each particle include two BOOs ($Q_{6}$ and $Q_{4}$), Voronoi cell volume, and C.N. This FCNN architecture mirrors the final component of the GNN (see Fig. 1). Similar to the GNN, we employed PyTorch with the Adam optimization algorithm, employing a learning rate of $10^{-6}$ and a batch size of 128. The evolution of the epoch-dependent loss function for both the training and test datasets is depicted in Supplementary Figs. S6-S9. Similar to Fig. 2(a), Fig. 4 depicts the fraction of accurate predictions by our FCNN for classification between two temperatures, $T_{1}$ and $T_{2}$, using two BOOs, $Q_{6}$ (a) and $Q_{4}$ (b), Voronoi cell volume (c), and C.N. (d). The results of AUC are shown in Fig. S10(a)-(d) of the supplementary material.
When $Q_{6}$ is used as the input variable, the structure at the lowest temperature, $T_{1}=0.21$, is clearly distinguished from those at other temperatures. However, distinguishing structures between high temperatures proves challenging, in contrast to the GNN results. When other structural descriptors, $Q_{4}$, Voronoi cell volume, and C.N., are used as input variables, the predictive performance follows a similar trend to that of $Q_{6}$, albeit with an overall degradation. These reductions in performance, particularly in the high-temperature regime, suggest that using physically-defined structural descriptors as input variables in the FCNN presents certain limitations. In particular, the results suggest that FCNNs with structural descriptors are more sensitive to thermal fluctuations at high temperatures. In contrast, the decrease in prediction accuracy for GNNs is relatively modest, implying that the GNN effectively suppresses the influence of thermal fluctuations and generates more robust features through the message passing process.
III.4 Attention coefficient and comparison with structural descriptors
The self-attention layer of our GNN architecture quantifies the normalized attention coefficient $\alpha_{vu}$ (see Fig. 1). As outlined in the Introduction, this coefficient $\alpha_{vu}$ evaluates the importance of node $u$’s feature to node $v$ in the graph structure. Specifically, strong relationships between nodes are reflected in larger attention coefficients, offering insights into the inter-particle correlations within amorphous structures. Here, the normalized attention coefficient, $\alpha_{vu}$, multiplied by the number of neighboring nodes of $v$, including node $v$ itself, is analyzed as the coordination-weighted attention coefficient, $\tilde{\alpha}_{vu}$ (see Eq. (8) in “Mehtods” for the definition of $\alpha_{vu}$ using the Softmax function). This multiplication accounts for the variations in the number of neighboring nodes for each node. Thus, a value of unity for $\tilde{\alpha}_{vu}$ represents the average level of the normalized attention coefficient, $\alpha_{vu}$.
It is essential to note that the attention coefficient solely serves as the foundation basis for the GNN predictions, even though it currently lacks explicit physical interpretation. Therefore, a comparative analysis between the attention coefficient and structural descriptors including BOOs ($Q_{6}$ and $Q_{4}$), Voronoi cell volume, and C.N., is significant. This analysis aims to provide deeper physical understanding into the high predictive performance of the GNN. To achieve this, we examine the correlations between structural descriptors of a given particle (node $v$) of interest and its edge-connected neighboring particle (node $u$), characterized by the corresponding coordination-weighted attention coefficient, $\tilde{\alpha}_{vu}$.
Figure 5 presents two-dimensional probability distributions for two BOOs, $Q_{6}$ (a), $Q_{4}$ (b), Voronoi cell volume, and C.N. (d) for the two temperatures, $T_{1}=0.21$ and $T_{2}=1.0$. These are averaged over 200 configurations (100 configurations from $T_{1}$ and 100 configurations of $T_{2}$ from the test dataset). The horizontal axis represents values for a given central particle (node $v$), while the vertical axis corresponds to values for its neighboring particle (node $u$). The distributions are further colored based on the corresponding $\tilde{\alpha}_{vu}$, averaged over the number of particle pairs stored in each two-dimensional bin. $\tilde{\alpha}_{vu}=1$ is also highlighted in red.
Note that the shape of the distribution is also symmetric, reflecting the inherent symmetry of structural descriptors between two particles. However, due to the asymmetry, $\tilde{\alpha}_{vu}$ and $\tilde{\alpha}_{uv}$, with respect to two nodes $v$ and $u$, as described in “Methods”, the color map exhibits asymmetry when considering both $\tilde{\alpha}_{vu}$ and $\tilde{\alpha}_{uv}$ by exchanging the roles of the tow particles. The results of other temperature combinations are shown in Figs. S11-S14 of the supplementary material.
Remarkably, Fig. 5(a) illustrates that neighboring particles with small $Q_{6}$ exhibit larger values ($\tilde{\alpha}_{vu}>1$). This finding suggests that neighboring particles with more disordered orientation play a significant role in the local environment of the central particle. A similar trend is observed for $Q_{4}$, though the contrast is less pronounced compared to $Q_{6}$, as shown in Fig. 5(b). In contrast, correlations with Voronoi cell volume are weaker (see Fig. 5(c)). Furthermore, Fig. 5(d) reveals that neighboring particles with large C.N. values also exhibit larger values ($\tilde{\alpha}_{vu}>1$). This trend suggests that, similar to $Q_{6}$, central particles tend to focus on disordered orientations in their neighboring environment.
These results demonstrates that the GNN effectively identifies static heterogeneities between locally disordered structures, enabling it to distinguish between the two temperatures. In particular, the correlation between coordination-weighted attention coefficients and structural descriptors, particularly $Q_{6}$ and C.N., arises from how the relative coordinates of neighboring particles for a given central particle are encoded into the graph data utilized by the GNN.
IV Conclusions
In this study, we utilized ML, specifically GNNs, to tackle the problem of classifying structural configurations generated by MD simulations of three-dimensional glass-forming liquids at various temperatures. In contrast to previous studies, [13, 28, 39] which utilize ML to differentiate between vitrified and liquid structures at a constant cooling rate, our study developed a GNN model capable of distinguishing between two different temperatures within a branch of the supercooled state.
The GNNs have yielded insights into the classification of structural configurations in glass-forming liquids by varying temperatures. Notably, we found that the structure of different temperatures with significantly different $\alpha$-relaxation times can be perfectly distinguished, while the ability to discriminate between states with close potential energies was degraded. This result suggests a convincing relationship between structural characteristics and slow dynamics observed in glass-forming liquids.
The adoption of GNNs proved highly advantageous, given their inherent capability to process graph data efficiently. By representing individual particles as nodes and their connections as edges within our graph-based framework, we utilized the advantages of GNNs to automatically generate feature representations, without relying on prior physical information. These features captured intricate relationships with neighboring particles and allowed us to discern distinct temperature-dependent structures. Notably, node features generated by the GNN exhibit correlations with physically-defined structural descriptors including the BOOs, $Q_{6}$ and $Q_{4}$, Voronoi cell volume, and C.N. The results demonstrates that the GNN learns information related to the local structure of the particle configuration from the amorphous structure.
Furthermore, we conducted a comparative analysis by applying conventional FCNNs to the same binary classification problem. Using $Q_{6}$, $Q_{4}$, Voronoi cell volume, and C.N. as input variables, this approach aimed to provide deeper insights into the fundamental mechanisms underpinning ML in glass-forming liquids. This analysis not only corroborated the significance of structural descriptors but also highlighted the distinct strengths of GNNs in capturing important features from amorphous configurations.
Finally, our investigation unveiled the significance of incorporating the self-attention mechanism into the GNN architecture, shedding light on the inter-node relationships within the system. This self-attention mechanism, originally introduced in Graph Attention Networks, enabled us to assess the importance of each node’s connections with its neighbors, quantifying the features associated with each node. This human-interpretable model was crucial in understanding the structural changes occurring with temperature variations. Swanson et al. reported that, in two-dimensional systems, edges with large attention coefficients tend to form a large connected graph in the glass, while several smaller disconnected graphs are observed in the liquid state [13]. By contrast, we provide physical interpretations to the coordination-weighted attention coefficient, $\tilde{\alpha}_{vu}$. Notably, our results reveal that small $Q_{6}$ values and large C.N. values of neighboring particles emerge as key features, both indicative of locally disordered structures. These findings emphasize the need for further analysis to deepen our understanding of the compelling relationship between static structural features and the dynamic behaviors of individual particles, aligning with the direction of studies by Bapst et al. and others [3].
Our exploration of GNNs for structural classification in glass-forming liquids opens several promising aspects for future research. Extending this work to encompass a broader range of materials, including different glass formers such as the widely used Kob–Andersen model, silica glasses, polymer glasses and supercooled water, will contribute to a more comprehensive understanding of the relationships between structural features and dynamics. In particular, the integration of additional structural descriptors parameters [70, 71], in conjunction with the GNN+self-attention framework, will valuable insights into the underlying mechanisms in glass- forming liquids. For silica glasses and supercooled water, combining this approach with structural descriptors, which captures the degree of local tetrahedral ordering, may provide deeper insights into the structural origins of their anomalous behaviors. [72] In addition, incorporating persistent homology could further advance the analysis from a topological perspective, even in polymer glasses lacking a well-defined structural motif. [73]
SUPPLEMENTARY MATERIAL
The supplementary material includes temperature dependence of $g(r)$, $F_{\mathrm{s}}(k,t)$, $\tau_{\alpha}$, and potential energy (Fig. S1); probability distribution of structural descriptors, including the BOOs, Voronoi cell volume, and C.N. (Fig. S2); epoch-dependent loss function of GNNs (Fig. S3); fraction of GNN accuracy and AUC for the temperature range 0.21-0.25 (Fig. S4); epoch-dependent loss function of GNNs for the temperature range 0.21-0.25 (Fig. S5); epoch-dependent loss function of FNCCs using structural descriptors (Fig. S6, S7, S8, and S9); fraction of FCNN AUC (Fig. S10); Probability distribution plots using structural descriptors for two temperature combination, shown alongsidethe coordination-weighted attention coefficient, $\tilde{\alpha}_{vu}$, for comparison (Fig. S11, S12, S13, and S14).
Acknowledgements.
We express our gratitude to H. Shiba for generously providing the information regarding the PyG_BOTAN code [74]. We also extend our thanks to N. Oyama for opening the code used in Ref. 28. We acknowledge T. Kawasaki and H. Mizuno for their insightful discussions. This work was supported by JSPS KAKENHI Grant-in-Aid Grant Nos. JP25K00968, JP24H01719, JP22H04542, JP22K03550, and JP23H02622. We acknowledge support from the Fugaku Supercomputing Project (Nos. JPMXP1020230325 and JPMXP1020230327) and the Data-Driven Material Research Project (No. JPMXP1122714694) from the Ministry of Education, Culture, Sports, Science, and Technology. The numerical calculations were performed at Research Center of Computational Science, Okazaki Research Facilities, National Institutes of Natural Sciences (Projects: 24-IMS-C051 and 25-IMS-C052) and at the D3 Center, The University of Osaka.
AUTHOR DECLARATIONS
Conflict of Interest
The authors have no conflicts to disclose.
Data availability statement
The data that support the findings of this study are openly available in Zenodo at https://doi.org/10.5281/zenodo.15151938. Further data that support the findings of this study are available from the corresponding author upon request.